John invested a part of $50000 at the bank which has 4 1/2% interest rate and the other part at the bank which has 5 1/2% The former yields $ 750 more each year than the latter. How much did he invest at ezch bank with different interest rate?
Answer by checkley71(8403)

.045X=.055(50000-X)+750
.045=2750-.055X+750
.045X+.055X=2750+750
.1X=3500
X=3500/.1
X=35000 AMOUNT INVESTED @ 4.5%.
50000-35000=15000 INVESTED @ 5.5%
PROOF
.045*35000=.055*15000+750
1575=825+750
1575=1575